# Exploiting Distribution Constraints for Scalable and Efficient Image Retrieval

### v1

@@ 1 Introduction
Image retrieval involves finding the closest match of a given image (query) in a vast database of images (often called the reference set). It has numerous applications, from product recommendations in e-commerce to place recognition in robotics. In general, state-of-the-art (SOTA) image retrieval systems are dataset-specific. They are trained in a contrastive manner on a training split of the data, typically hand-labeled for positive and negative pairs for each query. Imagine a scenario where a server has a vast reference database of fashion clothing images and receives query images from users to find the closest matches in the reference set. Creating a separate split and hand-labeling positive and negative pairs to train a dataset-specific model is infeasible. A more scalable solution is to use off-the-shelf feature extractors like Dino (Caron et al., 2021), DinoV2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and ViT (Dosovitskiy et al., 2021). However, the performance of off-the-shelf feature extractors is generally lower compared to dataset-specific models. Q1 (Scalability): Is there a way to enhance the performance of these off-the-shelf models in a completely unsupervised way by using just the reference set?
@@ 2 Related Work
Deep Metric Learning: Image retrieval is modeled as a Deep Metric Learning (DML) problem (Wang et al., 2017). DML focuses on learning latent representations where similar items are close and dissimilar items are far apart. Various techniques have been proposed (Oh Song et al., 2016; Schroff et al., 2015; Ustinova & Lempitsky, 2016; Sohn, 2016; Wang et al., 2019) to learn meaningful latent representations for effective retrieval, including contrastive loss (Hadsell et al., 2006) and triplet loss (Schroff et al., 2015). More recent works use advanced N-pair loss (Sohn, 2016) or multi-similarity loss (Wang et al., 2019). Analogous to image retrieval, Visual Place Recognition (VPR) is also modeled as a DML problem (Garg et al., 2021). Various training objectives have been proposed for VPR (Ge et al., 2020; Xiao et al., 2023; Leyva-Vallina et al., 2023; Berton et al., 2022) to learn latent representations for effective retrieval. However, all of these approaches are dataset-specific, meaning they train specific neural networks for each dataset.
Foundation Models in Image Retrieval: The rise of foundation models, such as DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), ViT (Dosovitskiy et al., 2021), has allowed recent works to explore the applicability of these off-the-shelf models in image retrieval. As shown in (Keetha et al., 2023; Lu et al., 2024), off-the-shelf models, in some cases, can have comparable retrieval performance to dataset-specific models. However, these approaches either use the off-the-shelf models as-is (Keetha et al., 2023) or adapt them using a labeled held-out dataset (Lu et al., 2024). In stark contrast, our approach, AE-SVC, exploits the inherent properties of the distribution to improve the performance of these off-the-shelf models without requiring any labeled data.
@@ 5.2 Results
Fig. 4 shows the performance comparison between AE-SVC and PCA across $4$ datasets, InShop, Stanford Online Products (SOP), Pittsburgh30K, and TokyoVal, using $3$ foundation models (ViT, DINOv2-Large, and DINO-B16). The performance of PCA (dotted line) represents the default off-the-shelf performance of the foundation models, while the performance of AE-SVC is shown with solid lines. AE-SVC consistently outperforms PCA across all datasets and embeddings. For example, the DINOv2-Large embedding (yellow) shows a $24\%$ improvement on InShop, $10\%$ on SoP, $5\%$ on Pittsburgh30K, and $22\%$ on TokyoVal at full embedding size. Overall, AE-SVC achieves an average improvement of $15.5\%$ across all datasets and embeddings at full size. Also, AE-SVC consistently surpasses PCA at smaller embedding sizes, providing a better size-performance trade-off. For additional results with more foundation models, see Appendix A.3.
Fig. 5 illustrates the advantages of (SS)2D on two datasets: InShop and Pittsburgh30K, utilizing two foundation models, Dinov2-Large and Dino-S8. We compare (SS)2D with a non-linear dimensionality reduction technique, Variational Auto Encoder (VAE) (black crosses). Additionally, we compare it with Similarity Space Distillation (SSD), which represents the theoretical upper bound of (SS)2D. Note that unlike VAE (black crosses) and SSD (red crosses), which are trained for each dimension separately, (SS)2D learns an adaptive embedding in one shot. From the plot, we observe that AE-SVC (orange) initially enhances the performance of the original embedding (blue), and (SS)2D subsequently provides additional improvements of up to $10\%$ at smaller embedding sizes. Also note that since SSD is trained for each dimension separately, it serves as the theoretical upper bound of (SS)2D. The results indicate that (SS)2D closely approaches this upper bound. For additional results involving more datasets and foundation models, please refer to Appendix A.4. As discussed in Sec. 1, a smaller embedding size directly translates to faster retrieval speeds. Therefore, the effective dimensionality reduction achieved with (SS)2D results in improved retrieval speeds, as demonstrated in Fig. 1(D).
@@ 6 Conclusion and Future Work
In this paper, we proposed AE-SVC and (SS)2D, which improve retrieval scalability and efficiency. Our results show a significant improvement on $4$ datasets, from image retrieval to place recognition tasks. AE-SVC demonstrates up to a $16\%$ improvement in retrieval performance, and (SS)2D shows an improvement of $10\%$ for smaller embedding sizes, which further boosts retrieval efficiency. Future work includes new retrieval models that orchestrate the variance-aware property in training loss of foundation models. Although we observe the resulting difference in the cosine similarity distributions of the foundation models and the data set-specific models (Figure 3), the fundamental discrepancy between the embedding spaces remains unknown to the community.
@@
In the theoretical analysis section (Sec. 4) of this work, we clearly state all the assumptions as and when necessary. Appendix A.1 and A.2 provides additional details necessary to understand the proofs in Sec. 4. We also discussed all the hyperparameters in Appendix A.6. Lastly, we release a reproducible codebase in the supplementary details. The code includes dataloaders, execution code, and links to download all the datasets and models used. We also provide detailed instructions on how to replicate our experiment setup and results, ensuring complete transparency and reproducibility.
@@ A.3 AE-SVC Additional Results with more foundation models
In Fig. 4, we showed the advantage of using AE-SVC for embeddings of three models. Here, in Fig. 6, we show additional results with three more embeddings, namely: CLIP, DINO-V2-Small, and DINO-S8. We observe similar trends as previously noted. We release our codebase in the supplementary material for easy reproduction of all these results.
@@
In Fig. 5, we showed the advantage of using (SS)2D on top of AE-SVC for additional improvement in retrieval performance at smaller dimensions. Here, in Fig. 7, we show additional results for (SS)2D on SOP with DINO-B16 and on TokyoVal with ViT. We observe similar trends as previously noted. For trying (SS)2D with more combinations of datasets and foundation models, please refer to our codebase that we release in the supplementary material.
Figure 7: Applying (SS)2D over AE-SVC leads to further performance boost at lower embedding sizes. Here, we see a similar trend as discussed in the main paper but with two additional foundation models and datasets.
@@ A.5 AE-SVC Additional Results on the Recall Metric
In Fig. 4, we show the advantage of using AE-SVC for embeddings of three models on the mean Average Precision at $\mathbf{k}$ (mAP@$\mathbf{k}$) metric. Here, in Fig. 8, we show additional results on the Recall@$\mathbf{k}$ metric. We observe very similar trends on the Recall@$\mathbf{k}$ metric as we did on the mAP@$\mathbf{k}$ metric. We release a reproducible codebase for readers to experiment with both metrics.

### v2

@@ 1 Introduction
Image retrieval involves finding the closest match of a given image (query) in a vast database of images (often called the reference set). It has numerous applications, from product recommendations in e-commerce to place recognition in robotics. In general, state-of-the-art (SOTA) image retrieval systems are dataset-specific. They are trained in a contrastive manner on a training split of the data, typically hand-labeled for positive and negative pairs for each query. Imagine a scenario where a server has a vast reference database of fashion clothing images and receives query images from users to find the closest matches in the reference set. Creating a separate split and hand-labeling positive and negative pairs to train a dataset-specific model is infeasible. A more scalable solution is to use off-the-shelf feature extractors like Dino (Caron et al., 2021), DinoV2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and ViT (Dosovitskiy et al., 2021). However, the performance of off-the-shelf feature extractors is generally lower compared to dataset-specific models. Q1 (Scalability): Is there a way to enhance the performance of these off-the-shelf models in a completely unsupervised way?
@@ 2 Related Work
Deep Metric Learning: Image retrieval is modeled as a Deep Metric Learning (DML) problem (Wang et al., 2017). DML focuses on learning latent representations where similar items are close and dissimilar items are far apart. Various techniques have been proposed (Oh Song et al., 2016; Schroff et al., 2015; Ustinova & Lempitsky, 2016; Sohn, 2016; Wang et al., 2019) to learn meaningful latent representations for effective retrieval, including contrastive loss (Hadsell et al., 2006) and triplet loss (Schroff et al., 2015). More recent works use advanced N-pair loss (Sohn, 2016) or multi-similarity loss (Wang et al., 2019). Analogous to image retrieval, Visual Place Recognition (VPR) is also modeled as a DML problem (Garg et al., 2021). Various training objectives have been proposed for VPR (Ge et al., 2020; Xiao et al., 2023; Leyva-Vallina et al., 2023; Berton et al., 2022; Shubodh et al., 2024) to learn latent representations for effective retrieval. However, all of these approaches are dataset-specific, meaning they train specific neural networks for each dataset.
Foundation Models in Image Retrieval: The rise of foundation models, such as DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), ViT (Dosovitskiy et al., 2021), has allowed recent works to explore the applicability of these off-the-shelf models in image retrieval. As shown in (Keetha et al., 2023; Lu et al., 2024), off-the-shelf models, in some cases, can have comparable retrieval performance to dataset-specific models. However, these approaches either use the off-the-shelf models as-is (Keetha et al., 2023; Omama et al., 2023; han Li et al., 2024) or adapt them using a labeled held-out dataset (Lu et al., 2024). In stark contrast, our approach, AE-SVC, exploits the inherent properties of the distribution to improve the performance of these off-the-shelf models without requiring any labeled data.
@@ 5.2 Results
Fig. 4 shows the performance comparison between AE-SVC and PCA across $4$ datasets, InShop, Stanford Online Products (SOP), Pittsburgh30K, and TokyoVal, using $3$ foundation models (ViT, DINOv2-Large, and DINO-B16). The performance of PCA (dotted line) represents the default off-the-shelf performance of the foundation models, while the performance of AE-SVC is shown with solid lines. AE-SVC consistently outperforms PCA across all datasets and embeddings. For example, the DINOv2-Large embedding (yellow) shows a $24\%$ improvement on InShop, $10\%$ on SoP, $2\%$ on Pittsburgh30K, and $22\%$ on TokyoVal at full embedding size. Overall, AE-SVC achieves an average improvement of $15.5\%$ across all datasets and embeddings at full size. Also, AE-SVC consistently surpasses PCA at smaller embedding sizes, providing a better size-performance trade-off. For additional results with more foundation models, see Appendix A.3.
Fig. 5 illustrates the advantages of (SS)2D on two datasets: InShop and TokyoVal, utilizing two foundation models, DINO-B16 and ViT. We compare (SS)2D with PCA (blue dashes) and a non-linear dimensionality reduction technique, Variational Auto Encoder (VAE) (black crosses). Additionally, we compare it with Similarity Space Distillation (SSD), which represents the theoretical upper bound of (SS)2D. Note that unlike VAE (black crosses) and SSD (red crosses), which are trained for each dimension separately, (SS)2D learns an adaptive embedding in one shot. From the plot, we observe that AE-SVC (orange) initially enhances the performance of the original embedding (blue), and (SS)2D subsequently provides additional improvements of up to $10\%$ at smaller embedding sizes. Also note that since SSD is trained for each dimension separately, it serves as the theoretical upper bound of (SS)2D. The results indicate that (SS)2D closely approaches this upper bound. For additional results involving more datasets and foundation models, please refer to Appendix A.4. As discussed in Sec. 1, a smaller embedding size directly translates to faster retrieval speeds. Therefore, the effective dimensionality reduction achieved with (SS)2D results in improved retrieval speeds, as demonstrated in Fig. 1(D).
@@ 6 Conclusion and Future Work
In this paper, we proposed AE-SVC and (SS)2D, which improve retrieval scalability and efficiency. Our results show a significant improvement on $4$ datasets, from image retrieval to place recognition tasks. AE-SVC demonstrates up to a $15.5\%$ improvement in retrieval performance, and (SS)2D shows an improvement of $10\%$ for smaller embedding sizes, which further boosts retrieval efficiency. Future work includes new retrieval models that orchestrate the variance-aware property in training loss of foundation models. Although we observe the resulting difference in the cosine similarity distributions of the foundation models and the data set-specific models (Figure 3), the fundamental discrepancy between the embedding spaces remains unknown to the community.
@@
In the theoretical analysis section (Sec. 4) of this work, we clearly state all the assumptions as and when necessary. Appendix A.1 and A.2 provides additional details necessary to understand the proofs in Sec. 4. We also discussed all the hyperparameters in Appendix A.6.
Acknowledgement
This work was supported in part by National Science Foundation grants No. 2133481 and No. 2148186, and by the Office of Naval Research (ONR) under Grant No. N00014-22-1-2254. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of these agencies.
@@ A.3 AE-SVC Additional Results with more foundation models
In Fig. 4, we showed the advantage of using AE-SVC for embeddings of three models. Here, in Fig. 6, we show additional results with three more embeddings, namely: CLIP, DINO-V2-Small, and DINO-S8. We observe similar trends as previously noted.
@@
In Fig. 5, we showed the advantage of using (SS)2D on top of AE-SVC for additional improvement in retrieval performance at smaller dimensions. Here, in Fig. 7, we show additional results for (SS)2D on more datasets and embedding combinations. We observe similar trends as previously noted.
Figure 7: Applying (SS)2D over AE-SVC leads to further performance boost at lower embedding sizes. Here, we see a similar trend as discussed in the main paper but with additional foundation models and datasets.
@@ A.5 AE-SVC Additional Results on the Recall Metric
In Fig. 4, we show the advantage of using AE-SVC for embeddings of three models on the mean Average Precision at $\mathbf{k}$ (mAP@$\mathbf{k}$) metric. Here, in Fig. 8, we show additional results on the Recall@$\mathbf{k}$ metric. We observe very similar trends on the Recall@$\mathbf{k}$ metric as we did on the mAP@$\mathbf{k}$ metric.
@@ A.6 Hyperparameters
A.7 Network Size Ablations
Table 1: Retrieval performance (map@4) with different sizes of AE-SVC models at varying embedding dimensions. As the network size increases, retrieval performance scales up. However, the results show diminishing returns and possible overfitting with deeper networks.
Model 8 32 64 128 256 384
AE-SVC1 0.04 0.25 0.332 0.388 0.403 0.413
AE-SVC2 0.093 0.340 0.450 0.510 0.520 0.550
AE-SVC3 0.093 0.342 0.461 0.538 0.556 0.568
AE-SVC5 0.093 0.340 0.400 0.450 0.480 0.432
AE-SVC10 0.01 0.06 0.09 0.10 0.17 0.20
PCA 0.04 0.250 0.324 0.370 0.403 0.411
[…]
Larger map@4 values ($\uparrow$) indicate better retrieval performance. The performance of linear AE-SVC (AE-SVC1) is comparable to PCA. As the network size increases, retrieval performance scales up, with AE-SVC3 achieving the best results across all dimensions. We have used AE-SVC3 in all our experiments across datasets and embedding types.
However, performance begins to drop at AE-SVC5 and significantly deteriorates at AE-SVC10, indicating diminishing returns and possible overfitting with deeper networks. This behavior is consistent across all embedding dimensions.
A.8 Constraint Losses Applied at Different Layers
Table 2 presents the retrieval performance (map@4) of various AE-SVC configurations across different embedding dimensions (8, 32, 64, 128, 256, and 384). This evaluation uses a 3-layer encoder-decoder architecture. The constraint losses, which consist of variance, covariance, and mean-centering losses, are applied at different stages of the network. AE-SVC applies these losses immediately after the encoder at the representation level. AE-SVC1, AE-SVC2, and AE-SVC3 apply these losses after the first, second, and third decoder layers, respectively. Since the network has a 3-layer encoder-decoder architecture, AE-SVC3 corresponds to applying these losses after the final decoder layer which aligns with existing representation learning methods such as Zbontar et al. (2021), Chen et al. (2020), and Kalantidis et al. (2021). In all configurations, a reconstruction loss is applied after the final decoder layer.
The proposed configuration (AE-SVC) consistently achieves the best retrieval performance across all embedding dimensions, highlighting the effectiveness of applying the loss at the representation level rather than later in the decoder network (like in previous approaches). Performance generally deteriorates when the loss is applied at deeper decoder layers, as observed in the results for AE-SVC1, AE-SVC2, and AE-SVC3.
Table 2: Retrieval performance (map@4) with constraints applied at different layer of the AE-SVC decoder. Consistent with our theoretical analysis in the main paper, the constraints should be applied directly after the encoder and we see a performance drop as they are applied later in the decoder.
Model 8 32 64 128 256 384
AE-SVC 0.093 0.342 0.461 0.538 0.556 0.568
AE-SVC1 0.090 0.342 0.444 0.518 0.538 0.550
AE-SVC2 0.010 0.340 0.436 0.473 0.480 0.484
AE-SVC3 0.090 0.330 0.404 0.443 0.463 0.471
A.9 Additional Results with More Baselines and Held-out Train Set
Figure 9: We show the retrieval performance of AE-SVC (orange) compared against PCA (blue), PCA Whitening (purple), TLDR (magenta) and Variational Autoencoder (black). For this experiment, we use the Dino-S8 feature extractor. While PCA Whitening and TLDR improve the performance of the default feature extractor, AE-SVC achieves a significantly greater improvement.
We have incorporated PCA Whitening (Hyvärinen et al., 2009) as a baseline for our approach (AE-SVC). We have also incorporated TLDR (Kalantidis et al., 2021), a recent dimensionality reduction work focussed on image retrieval. TLDR also focuses on dimensionality reduction for retrieval. However, TLDR uses only covariance loss. In stark contrast, we use a combination of three losses, including variance loss, which, as per our theoretical analysis, is the most important. Furthermore, following Barlow Twins, TLDR applies the covariance loss at the output of the decoder, whereas we apply it at the representation level.